OpenStudy (anonymous):

3. The linear equation represents an estimate of the average cost of gas for year x starting in 1997 (“Consumer price index,” 2006). The year 1997 would be represented by x = 1, for example, because it is the first year in the study. Similarly, 2005 would be year 9, or x = 9. a. What year would be represented by x = 4? b. What x-value represents the year 2018? c. What is the slope, or rate of change, of this equation? d. What is the y-intercept? e. What does the y-intercept represent? f. Assuming this growth trend continues, what will the price of gasoline be in the year 2018? How did

OpenStudy (anonymous):

a. is 2000

myininaya (myininaya):

how did you get that without an equation? I would just say these questions cannot be answered since there was not enough information given.

OpenStudy (anonymous):

im sorry the equation is y=0.5x+0.79

myininaya (myininaya):

so for the first one you plug in x=4 correct?

OpenStudy (anonymous):

after that im lost

myininaya (myininaya):

oh since x=4, then the year is 2000 ok gj now for b....

myininaya (myininaya):

so x=1 we have 1997 x=2 we have 1998 x=3 we have 1999 x=4 we have 2000 and so on.. so the equation for the year would be Z=1996+x lets see for x=1, we get Z=1996+1=1997 for x=2 we get Z=1996+2=1998 and so on... so the question is what is x if Z=2018 so we have 2018=1996+x solve this equation for x

myininaya (myininaya):

what did you get when you solved the equation that i just typed for x?

OpenStudy (anonymous):

x=22

myininaya (myininaya):

ok cool so did you understand everything for b?

OpenStudy (anonymous):

no not really by u breaking it down help me find the answer

myininaya (myininaya):

I just made up a variable for the year

OpenStudy (anonymous):

can we do the next one c

myininaya (myininaya):

k it wants to know the slope of the line? what is the slope of any line in this form y=mx+b?

OpenStudy (anonymous):

m=

OpenStudy (anonymous):

is that correct

myininaya (myininaya):

yes

myininaya (myininaya):

so y=.5x+79 has slope_________?

myininaya (myininaya):

what is m?

OpenStudy (anonymous):

ok 1 min

myininaya (myininaya):

slope=m=.5

myininaya (myininaya):

ok

OpenStudy (anonymous):

oh ok the slope is .5 and would b=79

OpenStudy (anonymous):

or the answer to the ? is just .5

myininaya (myininaya):

the question c is just looking for slope so it would be .5 the question d is just looking for the y intercept so it would be .79 (I forgot the decimal when I typed it a sec ago)

OpenStudy (anonymous):

therefore the y-intersect would b =0

myininaya (myininaya):

the y-intercept is when x=0

OpenStudy (anonymous):

than im lost on problem e

myininaya (myininaya):

ok we will come back to it then lets look at the last one average gas is y=.5x+.79 and since we are looking at yr 2018 then x=22 so the average gas for yr 2018 is y=.5(22)+.79

OpenStudy (anonymous):

y=11.79

myininaya (myininaya):

ok back to e. the y-intercept is when x=0 because if we were to graph a point on the y axis x would always be zero since we aren't going left and right any from the origin. for example graph (0,5) and (0,-6). where are these points?

OpenStudy (anonymous):

would it be -11